OpenStudy (anonymous):

she invested $5000 into a fund that is expected to grow by 3.5% per year. How long will it take the fund she invested in to be worth $10,000?

OpenStudy (whpalmer4):

This is compound growth, I assume?

OpenStudy (anonymous):

I'm assuming so , yes.

OpenStudy (whpalmer4):

Do you know the formula for calculating compound growth?

OpenStudy (anonymous):

i dont think so .

OpenStudy (whpalmer4):

If we have a growth rate of \(r\) (expressed as a decimal), after 1 growth period, our initial quantity will go from \(P\) to \(P*(1+r)\). After 2 growth periods, it will be \(P*(1+r)*(1+r) = P*(1+r)^2\) After 3 growth periods, it will be \(P*(1+r)*(1+r)*(1+r) = P*(1+r)^3\) After \(n\) growth periods, it will be \(P*(1+r)^n\)

OpenStudy (whpalmer4):

So, what you're trying to find is the value of \(n\) such that \[10000 = 5000*(1+0.035)^n\] Any idea how to proceed?

OpenStudy (anonymous):

solve for n .

OpenStudy (whpalmer4):

Well, yes, but do you know how to do so?

OpenStudy (anonymous):

No .

OpenStudy (whpalmer4):

First step would be to divide both sides by 5000 \[\frac{10000}{5000} = \frac{5000(1+0.035)^n}{5000}\]\[2 = (1+0.035)^n\] Do you know about logarithms?

OpenStudy (whpalmer4):

We'll use a property of logarithms that \[\log x^a = a\log x\] then take the log of both sides: \[\log 2 = \log (1.035)^n\]\[\log 2 = n\log 1.035\]\[n = \frac{\log 2}{\log 1.035}\] Now you break out your calculator and compute the value of \(n\)...

OpenStudy (whpalmer4):

I'll give you a hint: \(\log 2 \approx 0.30103\)

OpenStudy (whpalmer4):

Not really a hint, I guess, as much as saving you a step.